A uniform ladder AB of weight 180N and length 8m, rest in equilibrium with its foot on a rough horizontal ground and its upper end against a rough vertical wall. The ladder makes an angle of 60 degrees with the horizontal ground. The coefficient of friction between the ladder and the wall is 1/4 and that between the ladder and the ground is 1/2. Find the normal reactions at A and B
1
Expert's answer
2020-06-24T11:21:23-0400

"\theta = 60^\circ, \;\; \mu_1 = 0.5, \;\; \mu_2 = 0.25." The length "l = 8\,\mathrm{m}," the weight "mg = 180\,\mathrm{N}." Let "N_1" be the normal reaction at the foot of the ladder and "N_2" be the normal reaction at the highest point of the ladder. We may write two equations of the force balance in projection on vertical and horizontal axes:

"F_{fr,2} + N_1= mg\,; \;\; N_2 = F_{fr,1}\,."

Let us assume that the friction is proportional to the friction of rest, so "\mu_2N_2\nu + N_1 = mg, \;\; N_2=\mu_1N_1\nu."

We may also write equation concerning torques using the top of ladder:

at the top: "mg\dfrac l2\cos\alpha = N_1l\cos\alpha - F_{fr,1}l\sin\alpha."

Substituting "F_{fr,1}\,, \; F_{fr,2}" from the balance of forces, we get

"mg\dfrac 12\cos\alpha =N_1\cos\alpha -N_2\sin\alpha" . Therefore, "N_1 = \dfrac{mg}{2} + N_2\tan\alpha."

So we obtain the system of equations

"\begin{cases}
\mu_2N_2\nu + N_1 = mg, \\ N_2=\mu_1N_1\nu, \\
N_1 = \dfrac{mg}{2} + N_2\tan\alpha.
\end{cases}" "\begin{cases}
\mu_1\mu_2N_1\nu^2 + N_1 = mg, \\ N_2=\mu_1N_1\nu, \\
N_1 = \dfrac{mg}{2} + \mu_1N_1\nu\tan\alpha.
\end{cases}"

Then we get equation "\mu_1\mu_2\nu^2+2\mu_1\nu\tan\alpha-1=0, \;\;" "\nu\approx 0.555."

"N_1 = \dfrac{mg}{\mu_1\mu_2\nu^2+1} = 173\,\mathrm{N}."

"N_2 = 0.5\cdot N_2\cdot 0.555 = 48\,\mathrm{N}."